I would like help with this problem:
1/2q - 6 = 1/5q
What I have worked so far is:
1/2q - 6 - 1/5q = 1/5q - 1/5q
Could someone help me from here?

Answer by tutorcecilia(2152)

1/2q - 6 - 1/5q = 1/5q - 1/5q
1/2q - 1/5q - 6 = 0
3/10q - 6 = 0
3/10q - 6 + 6 = 0 + 6
3/10q = 6
(3/10q)(10/3) = 6(10/3)
q = 20
.
Check by plugging q=20 back into the original equation:
1/2(20) - 6 = 1/5(20)
10 - 6 = 4
4 = 4 Checks out.